### v1

# Extreme Score Distributions in Countable-Outcome Round-Robin Tournaments of Equally Strong Players

Abstract
We consider a general class of round-robin tournament models of equally strong players. In these models, each of the $n$ players competes against every other player exactly once. For each match between two players, the outcome is a value from a countable subset of the unit interval, and the scores of the two players in a match sum to one. The final score of each player is defined as the sum of the scores obtained in matches against all other players. We study the distribution of extreme scores, including the maximum, second maximum, and lower-order extremes. Since the exact distribution is computationally intractable even for small values of $n$, we derive asymptotic results as the number of players $n$ tends to infinity, including limiting distributions, and rates of convergence.
Keywords: Complete graph, negative association, order statistics, Poisson approximation, total variation distance
MSC2020: 62G32; 05C20; 60C05
1 Introduction and Background
Round-robin tournaments constitute a classical and extensively studied model in combinatorics, probability, and statistics. In such tournaments, each of the $n$ players competes against each of the others n1n-1 players, and the outcome of each match contributes to the players’ final scores. The structure and distribution of these scores have been studied from combinatorial, probabilistic, and statistical perspectives; however, the behavior of extreme values has received comparatively little attention.
In an 1895 paper, Edmund Landau proposed and investigated an alternative scoring system for chess tournaments, differing from the classical scheme—still in use today—in which the winner of a game receives $1$ point, the loser 0 points, and, in the case of a draw, each player receives $\tfrac{1}{2}$ point (Landau, 1895, 1914). More recently, Sinn and Ziegler (2022) observed that Landau’s idea of determining scores through relative comparisons plays a central role in the original 1998 Google search algorithm, PageRank.
Another eminent mathematician, Ernst Zermelo, referred to Landau’s 1914 work (Landau, 1914) and proposed a probabilistic model in which the probability that a player wins a game is given by the ratio of that player’s strength to the sum of the strengths of the two competing players. He estimated the model parameters using the method of maximum likelihood and applied this approach to the analysis of the 1924 New York chess tournament (Zermelo, 1929). The Zermelo model was later rediscovered by Bradley and Terry (Bradley and Terry, 1952) and by Ford (Ford, 1957); see also the detailed historical discussion in (David and Edwards, 2001), as well as the probabilistic aspects of rating models discussed in (Aldous, 2017).
Formally, in a round-robin tournament, each of the $n$ players competes against each of the other n1n-1 players exactly once. When player $i$ competes against player $j$, the reward obtained by player $i$ is a random variable $X_{ij}$. We define
$s_{i}(n)=\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}X_{ij}$
to be the score of player $i$, $1\leq i\leq n$, after playing all other n1n-1 players.
Throughout this paper, we distinguish between models that allow different sets of possible values for the random variables $X_{ij}$. Specifically, we consider the following cases:
  1. 1.
    $D_{k}=\left\{0,\frac{1}{k},\frac{2}{k},\ldots,1\right\}$,
  2. 2.
    $D\subset[0,1]$ is a countable set.
In the classical round-robin tournament, $X_{ij}+X_{ji}=1$ and $X_{ij}\in D_{1}$, whereas in the chess round-robin tournament discussed above, $X_{ij}+X_{ji}=1$ and $X_{ij}\in D_{2}$. We assume that the $\binom{n}{2}$ random vectors $\{(X_{ij},X_{ji}):1\leq i<j\leq n\}$ are independent.
We refer to the vector
$\bigl(s_{1}(n),s_{2}(n),\ldots,s_{n}(n)\bigr)$
as the score sequence of the tournament. The corresponding order statistics are denoted by
$\bigl(s_{(1)}(n)\leq s_{(2)}(n)\leq\cdots\leq s_{(n)}(n)\bigr).$
We also denote by
$\bigl(s_{1}^{\star}(n),s_{2}^{\star}(n),\ldots,s_{n}^{\star}(n)\bigr)$
the normalized score sequence of the tournament, where, for a given model, each score is centered by subtracting its expectation and scaled by dividing by its standard deviation. The corresponding order statistics are denoted by
$\bigl(s_{(1)}^{\star}(n)\leq s_{(2)}^{\star}(n)\leq\cdots\leq s_{(n)}^{\star}(n)\bigr).$
The exact distribution of score sequences in the classical round-robin tournament for $n\leq 9$, and in the chess round-robin model for $n\leq 6$ under the assumption of equally strong players, was determined by the prominent mathematician Percy A. MacMahon (MacMahon, 1923a, b). However, even for small values of $n$, determining the exact distribution is in general computationally infeasible due to the astronomical size of the sample space, which, for example, in the classical round-robin tournament equals $2^{\binom{n}{2}}$.
An excellent introduction to the combinatorial and probabilistic aspects of classical round-robin tournaments, together with references to earlier work, is provided in Moon (1968); see also the corrected 2013 version available via Project Gutenberg.
Round-robin tournaments may be viewed as collections of paired comparisons among players. The combinatorial, probabilistic, and statistical aspects of paired-comparison models—often referred to as the method of paired comparisons, a term coined by Thurstone (Thurstone, 1927); see also (David and Edwards, 2001)—have been extensively studied in the literature. Classical and modern contributions include, among many others, (Kendall and Babington Smith, 1940; Kendall, 1955; Bose, 1956; David, 1959; Huber, 1963; David, 1988; Joe, 1988; Chebotarev and Shamis, 1998; Alon et al., 2006; Adler et al., 2017; Chetrite et al., 2017; Bruss and Ferguson, 2018; Csató, 2021; Ross, 2022). Closely related problems in ranking and learning from pairwise comparisons have also attracted considerable attention; see, for example, (Heckel et al., 2019; Newman, 2023). A comprehensive treatment of the statistical aspects of paired comparisons, together with combinatorial methods, experimental designs, and selection and ranking procedures, is provided in the monograph by Herbert A. David (David, 1988).
The behavior of extreme scores in round-robin tournaments has received relatively little explicit attention in the literature, although the earliest result known to us appears already in (Huber, 1963). The investigation of such behavior is closely related to negative dependence among the scores, which we discuss next, together with relevant references, previous results, and a precise problem formulation.
We define the following notions of negative dependence. See Joag-Dev and Proschan (1983) and the references therein for details. Throughout this paper, the terms increasing and decreasing are understood to mean nondecreasing and nonincreasing, respectively.
Definition 1.
(Joag-Dev and Proschan (1983), Definition 2.3) The random variables $S_{1},\ldots,S_{n}$, or the vector ${\bf S}=(S_{1},\ldots,S_{n})$, are said to be negatively lower orthant dependent (NLOD) if, for all $s_{1},\ldots,s_{n}\in\mathbb{R}$,
$\mathbb{P}\left(S_{1}\leq s_{1},\ldots,S_{n}\leq s_{n}\right)\leq\mathbb{P}\left(S_{1}\leq s_{1}\right)\cdots\mathbb{P}\left(S_{n}\leq s_{n}\right),$ (1)
and negatively upper orthant dependent (NUOD) if
$\mathbb{P}\left(S_{1}>s_{1},\ldots,S_{n}>s_{n}\right)\leq\mathbb{P}\left(S_{1}>s_{1}\right)\cdots\mathbb{P}\left(S_{n}>s_{n}\right).$ (2)
Negative orthant dependence (NOD) is said to hold if both (1) and (2) hold.
The notion of negative association was studied early in Alam and Saxena (1981); we recall the standard definition in the form given in Joag-Dev and Proschan (1983).
Definition 2.
(Joag-Dev and Proschan (1983), Definition 2.1) The random variables $S_{1},\ldots,S_{n}$ or the vector ${\bf S}=(S_{1},\ldots,S_{n})$ are said to be negatively associated (NA) if for every pair of disjoint subsets $A_{1},A_{2}$ of $\left\{1,2,\ldots,n\right\}$,
$\mathrm{Cov}\left(f_{1}(S_{i},i\in A_{1}),f_{2}(S_{j},j\in A_{2})\right)\leq 0,$
whenever $f_{1}$ and $f_{2}$ are real-valued functions, increasing in all coordinates.
NA implies NOD (see Joag-Dev and Proschan (1983)).
For the classical round-robin tournament model with $p_{ij}=\mathbb{P}(X_{ij}=1)=\tfrac{1}{2}$, Huber (Huber, 1963) proved that
$s_{(n)}^{\star}(n)-\sqrt{2\log(n-1)}\xrightarrow{\mathbb{P}}0\quad\text{as }n\to\infty,$ (3)
where $\log(x)$ denotes the natural logarithm. One of the key elements in his proof was establishing the NLOD property (1) of the score vector $\bigl(s_{1}(n),s_{2}(n),\ldots,s_{n}(n)\bigr)$ via a coupling argument. For round-robin tournaments with repeated games, where $X_{ij}\sim\mathrm{Bin}(r_{ij},p_{ij})$ independently for all $i<j$ and $X_{ij}+X_{ji}=r_{ij}$, Ross (Ross, 2022) established the NLOD property (1) for the score vector $\bigl(s_{1}(n),s_{2}(n),\ldots,s_{n}(n)\bigr)$, which allows the study of probabilities of the form $\mathbb{P}\!\left(s_{i}(n)>\max_{j\neq i}s_{j}(n)\right)$. Under the model of Ross (2022), with integer-valued $X_{ij}$ but allowing for an arbitrary discrete distribution in place of the binomial one, Malinovsky and Moon (2022) established the NLOD property (1) for the score vector $\bigl(s_{1}(n),s_{2}(n),\ldots,s_{n}(n)\bigr)$ and used this property to derive a Huber-type result of the form (3). For the chess round-robin model with equally strong players, Malinovsky (2022a, b) obtained the asymptotic distributions of the maximal score $s_{(n)}(n)$, the second maximal score $s_{(n-1)}(n)$, and, more generally, of all order statistics down to $s_{(1)}(n)$. Finally, Epstein (1967) stated, without proof, that in a classical round-robin tournament with equally strong players, the probability that there is a unique winner with the maximal score tends to $1$ as $n\to\infty$. This conjecture was proved in Malinovsky and Moon (2024). More recently, Amir and Malinovsky (2026) extended this result to the case in which the match outcomes $X_{ij}$ take values in a countable set $D\subset[0,1]$.
It is interesting to note that related types of problems have been extensively studied in the theory of random graphs, where the nature of dependence is, in a sense, opposite to that arising in round-robin tournaments, which exhibit negative dependence. A round-robin tournament can be represented as a complete oriented graph on $n$ vertices, where the vertices correspond to players. For each unordered pair of distinct vertices $i$ and $j$, exactly one directed edge is present: it is oriented from $i$ to $j$ if $X_{ij}=a$, and from $j$ to $i$ if $X_{ji}=1-a$. By contrast, in the Erdős–Rényi random graph model Erdős and Rényi (1959, 1960), there are $n$ vertices, and each unordered pair of distinct vertices $i$ and $j$ is connected by an undirected edge independently with probability $p$ (which may depend on $n$). If an edge is present, no orientation is assigned. The seminal papers Erdős and Rényi (1959, 1961) initiated the study of the asymptotic behavior of extreme degrees (maximal and minimal) in random graphs. Subsequently, Ivchenko (1973) studied the asymptotic marginal distributions of the order statistics of vertex degrees in the sparse regime where $p=p(n)\to 0$ as $n\to\infty$, while Bollobás (Bollobás, 1980, 1981) independently established corresponding results for fixed $p$, with further generalizations given in (Bollobás, 2001); see also (Vakhrushev and Zhukovskii, 2024; Malinovsky, 2024).
In this paper, we extend previous results on the asymptotic distributions of the order statistics $s_{(i)}(n)$, $i=1,\ldots,n$, obtained in (Malinovsky, 2022a, b) for the chess round-robin tournament model, to a more general setting in which the match outcomes $X_{ij}$ take values in a countable set $D\subset[0,1]$. In addition, we establish rates of convergence.
The remainder of the paper is organized as follows. In Section 2 we introduce the general round-robin tournament model with countable outcomes and provide additional notation and technical facts. Section 3 presents the main results, including the asymptotic distributions of the order statistics and rates of convergence. Section 4 contains the proofs of the main results. Appendices A, B, C provide proofs of auxiliary results and necessary lemmas, theorems used in Section 4.
2 Model and Additional Notation
Model $M_{[0,1]}$. Let $D\subset[0,1]$ be a countable set of possible values of $X_{ij}$, and assume that $|D|>1$. We consider a round-robin tournament in which, for all $i\neq j$,
$X_{ij}+X_{ji}=1,\qquad X_{ij}\in D.$ (4)
In particular, $a\in D$ implies $1-a\in D$.
We assume that the players are equally strong and each value can be attained with positive probability, that is,
$0<\mathbb{P}\!\left(X_{ij}=a\right)=\mathbb{P}\!\left(X_{ji}=a\right),\qquad\text{for all }a\in D.$ (5)
Under (4), the assumption (5) is equivalent to the distribution of $X_{ij}$ being symmetric about $1/2$: for every $a\in D$,
$0<\mathbb{P}(X_{ij}=a)=\mathbb{P}(X_{ij}=1-a),\qquad 1\leq i<j\leq n.$ (6)
Consequently, the scores $s_{1}(n),\ldots,s_{n}(n)$ are identically distributed.
Under Model $M_{[0,1]}$, let
$\mu=\mathbb{E}(X_{ij}),\qquad\sigma=\sqrt{\mathrm{Var}(X_{ij})}.$
Let
$a_{n}=(2\log n)^{-1/2},\qquad b_{n}=(2\log n)^{1/2}-\frac{\log\log n+\log(4\pi)}{2(2\log n)^{1/2}},$ (7)
and for $t\in\mathbb{R}$ define
$x_{n}(t)=b_{n}+a_{n}t=(2\log n)^{1/2}+\frac{t-\frac{\log\log n+\log(4\pi)}{2}}{(2\log n)^{1/2}}.$ (8)
Under Model $M_{[0,1]}$, the normalized score of player $i$ is
$s^{\star}_{i}(n)=\frac{s_{i}(n)-(n-1)\mu}{\sqrt{n-1}\,\sigma}.$
Define the exceedance indicators and their count:
$I_{i}^{(n)}(t):=\mathbf{1}\{s^{\star}_{i}(n)>x_{n}(t)\},\qquad W_{n}(t):=\sum_{i=1}^{n}I_{i}^{(n)}(t).$ (9)
Write
$p_{n}(t):=\mathbb{P}(s^{\star}_{i}(n)>x_{n}(t)),\qquad\lambda_{n}(t):=\mathbb{E}[W_{n}(t)]=np_{n}(t).$ (10)
The total variation distance between the distributions of random elements $X$ and $Y$ is denoted $d_{\mathrm{TV}}(X,Y)$ and defined by
$d_{\mathrm{TV}}(X,Y)=\frac{1}{2}\sum_{k\geq 0}\left|\mathbb{P}(Y=k)-\mathbb{P}(X=k)\right|.$
Fix an integer $j\geq 0$ and define
$M_{n,j}:=\frac{s_{(n-j)}^{\star}(n)-b_{n}}{a_{n}}.$
The notation $f_{n}\sim g_{n}$ denotes asymptotic equivalence, that is, $\lim_{n\to\infty}\frac{f_{n}}{g_{n}}=1.$
3 Main Results
Theorem 1.
For every fixed $t\in\mathbb{R}$, there exists a constant $C(t)<\infty$ such that
$d_{\mathrm{TV}}\bigl(\mathcal{L}(W_{n}(t)),\,\mathrm{Poisson}(e^{-t})\bigr)\leq C(t)\,\frac{(\log\log n)^{2}}{\log n},\qquad n\to\infty.$ (11)
That is, $W_{n}(t)$ converges in total variation to a Poisson random variable with mean $\mathrm{e}^{-t}$, at a rate of order $(\log\log n)^{2}/\log n$. Convergence in total variation implies convergence in distribution (see, for example, (Grimmett and Stirzaker, 2020, Exercise 7.2.9, p. 362; Exercise 7.11.16, p. 401)), and therefore we obtain the following corollary.
Corollary 1.
For every fixed $t\in\mathbb{R}$ and each fixed integer $k\geq 0$,
$\lim_{n\to\infty}\mathbb{P}\!\left(W_{n}(t)=k\right)=\mathrm{e}^{-\mathrm{e}^{-t}}\frac{\mathrm{e}^{-tk}}{k!}.$
Moreover, the convergence of $W_{n}(t)$ to a Poisson random variable with mean $\mathrm{e}^{-t}$ holds in total variation at rate of order $(\log\log n)^{2}/\log n$.
Furthermore, recalling (9), we observe that
$\{W_{n}(t)\leq j\}=\{s^{\star}_{(n-j)}(n)\leq x_{n}(t)\}.$
Combining this observation with Corollary 1, we obtain the following corollary.
Corollary 2.
For every fixed $t\in\mathbb{R}$ and each fixed integer $j\geq 0$,
$\lim_{n\to\infty}\mathbb{P}\!\left(\frac{s_{(n-j)}(n)-(n-1)\mu}{\sqrt{n-1}\,\sigma}\leq x_{n}(t)\right)=\exp(-e^{-t})\sum_{k=0}^{j}\frac{e^{-tk}}{k!},$ (12)
where $x_{n}(t)$ is defined in (8). Moreover, the convergence in (12) holds at a rate of order $(\log\log n)^{2}/\log n$.
4 Proofs
Proof of Theorem 1.
Under Model $M_{[0,1]}$, it follows from (Malinovsky and Rinott, 2023, Proposition 1) that the score vector $\bigl(s_{1}(n),\ldots,s_{n}(n)\bigr)$ is negatively associated (see Definition 2). Property $P_{6}$ in Joag-Dev and Proschan (1983) states that increasing functions defined on disjoint subsets of a negatively associated family are themselves negatively associated.
For each $i=1,\ldots,n$, the indicator $I_{i}^{(n)}(t)=\mathbf{1}\{s_{i}^{\star}(n)>x_{n}(t)\}$ is an increasing function of the single coordinate $s_{i}(n)$ (since $s_{i}^{\star}(n)$ is obtained from $s_{i}(n)$ by centering and scaling). Hence, by Property $P_{6}$ in Joag-Dev and Proschan (1983), the family $\bigl(I_{1}^{(n)}(t),\ldots,I_{n}^{(n)}(t)\bigr)$ is negatively associated.
Using this fact, and combining Theorem 2.I (p. 30) and Corollary 2.C.2 (p. 36) of Barbour et al. (1992), we obtain for every fixed $t\in\mathbb{R}$ the following total variation bound:
$d_{\mathrm{TV}}\bigl(\mathcal{L}(W_{n}(t)),\,\mathrm{Poisson}(\lambda_{n}(t))\bigr)$ $=\frac{1}{2}\sum_{k\geq 0}\left|\mathbb{P}(W_{n}(t)=k)-\mathbb{P}(\mathrm{Poisson}(\lambda_{n}(t))=k)\right|$
$\leq\left(1-\mathrm{e}^{-\lambda_{n}(t)}\right)\left(1-\frac{\mathrm{Var}(W_{n}(t))}{\lambda_{n}(t)}\right)$
$=\left(1-\mathrm{e}^{-\lambda_{n}(t)}\right)\left[\frac{\lambda_{n}(t)}{n}-\frac{n(n-1)\,\mathrm{Cov}\!\left(I_{1}^{(n)}(t),I_{2}^{(n)}(t)\right)}{\lambda_{n}(t)}\right]$
$\leq\frac{\lambda_{n}(t)}{n}-\frac{n(n-1)\,\mathrm{Cov}\!\left(I_{1}^{(n)}(t),I_{2}^{(n)}(t)\right)}{\lambda_{n}(t)}.$ (13)
Here the equality follows from the fact that $s_{1}(n),\ldots,s_{n}(n)$ are identically distributed and from the definition of $\lambda_{n}(t)$ in (10).
To evaluate the upper bound, we will need the following two propositions.
Proposition 1.
For every fixed $t\in\mathbb{R}$,
$\lambda_{n}(t)=n\,p_{n}(t)=\mathrm{e}^{-t}\left(1+O\!\left(\frac{(\log\log n)^{2}}{\log n}\right)\right),\qquad n\to\infty.$
Proof.
Appendix B. ∎
Proposition 2.
For every fixed $t\in\mathbb{R}$,
$\mathrm{Cov}\!\left(I_{1}^{(n)}(t),I_{2}^{(n)}(t)\right)=-\frac{2\,\mathrm{e}^{-2t}\log n}{n^{3}}\left(1+O\!\left(\frac{(\log\log n)^{2}}{\log n}\right)\right),\qquad n\to\infty.$
Proof.
Appendix C. ∎
By the triangle inequality,
$d_{\mathrm{TV}}\bigl(\mathcal{L}(W_{n}(t)),\,\mathrm{Poisson}(e^{-t})\bigr)$ $\leq d_{\mathrm{TV}}\bigl(\mathcal{L}(W_{n}(t)),\,\mathrm{Poisson}(\lambda_{n}(t))\bigr)+d_{\mathrm{TV}}\bigl(\mathrm{Poisson}(\lambda_{n}(t)),\,\mathrm{Poisson}(e^{-t})\bigr).$ (14)
From Proposition 1 and Proposition 2, we obtain
$-\frac{n(n-1)\,\mathrm{Cov}\!\left(I_{1}^{(n)}(t),I_{2}^{(n)}(t)\right)}{\lambda_{n}(t)}$ $=\frac{n(n-1)}{\lambda_{n}(t)}\cdot\frac{2e^{-2t}\log n}{n^{3}}\left(1+O\!\left(\frac{(\log\log n)^{2}}{\log n}\right)\right)$
$=\frac{2e^{-2t}\log n}{\lambda_{n}(t)\,n}\left(1+O\!\left(\frac{(\log\log n)^{2}}{\log n}\right)\right)$
$=\frac{2e^{-t}\log n}{n}\left(1+O\!\left(\frac{(\log\log n)^{2}}{\log n}\right)\right).$ (15)
Appealing again to Proposition 1 and using (15), we obtain from (4) that for every fixed $t\in\mathbb{R}$ there exists a constant $C_{1}(t)<\infty$ such that
$d_{\mathrm{TV}}\bigl(\mathcal{L}(W_{n}(t)),\,\mathrm{Poisson}(\lambda_{n}(t))\bigr)\leq C_{1}(t)\,\frac{\log n}{n},\qquad n\to\infty.$ (16)
From Result 1 in Appendix A, for all $\lambda,\mu\geq 0$,
$d_{\mathrm{TV}}\bigl(\mathrm{Poisson}(\lambda),\,\mathrm{Poisson}(\mu)\bigr)\leq|\lambda-\mu|.$ (17)
Hence, by (17) and Proposition 1,
$d_{\mathrm{TV}}\bigl(\mathrm{Poisson}(\lambda_{n}(t)),\,\mathrm{Poisson}(e^{-t})\bigr)\leq|\lambda_{n}(t)-e^{-t}|=e^{-t}\,O\!\left(\frac{(\log\log n)^{2}}{\log n}\right),\qquad n\to\infty.$ (18)
Substituting (16) and (18) into (14) gives
$d_{\mathrm{TV}}\bigl(\mathcal{L}(W_{n}(t)),\,\mathrm{Poisson}(e^{-t})\bigr)$ $\leq C_{1}(t)\,\frac{\log n}{n}+e^{-t}\,O\!\left(\frac{(\log\log n)^{2}}{\log n}\right).$
Since $\frac{\log n}{n}=o\!\left(\frac{(\log\log n)^{2}}{\log n}\right)$, we obtain (11). ∎
Appendix A Auxiliary Results
Result 1.
For all $\lambda,\mu\geq 0$,
$d_{\mathrm{TV}}\bigl(\mathrm{Poisson}(\lambda),\,\mathrm{Poisson}(\mu)\bigr)\leq|\lambda-\mu|.$
Proof of Result 1.
Assume without loss of generality that $\lambda\leq\mu$. Couple $N\sim\mathrm{Poisson}(\lambda)$ and $M\sim\mathrm{Poisson}(\mu)$ by letting $M=N+K$, where $K\sim\mathrm{Poisson}(\mu-\lambda)$ is independent of $N$. Then
$d_{\mathrm{TV}}(N,M)\leq\mathbb{P}(M\neq N)=\mathbb{P}(K>0)=1-e^{-(\mu-\lambda)}\leq\mu-\lambda=|\lambda-\mu|,$
where the first inequality follows from (Barbour et al., 1992, equation (1.4), p. 254). ∎
Let $\Phi(t)=\int_{-\infty}^{t}\varphi(x),dx$, where $\varphi(x)=(2\pi)^{-1/2}e^{-x^{2}/2}$ denotes the density of the standard normal distribution.
Result 2 ((Feller, 1968, Chapter VII.7, p. 193)).
$1-\Phi(x)=\frac{\varphi(x)}{x}\left(1-\frac{1}{x^{2}}+O\!\left(\frac{1}{x^{4}}\right)\right),\qquad x\to\infty.$
Lemma 1.
Let $v_{n}\to\infty$ and let $u_{n}$ satisfy $u_{n}=v_{n}+\Delta_{n},\,\,\,\Delta_{n}=o(1).$
Then, as $n\to\infty$,
$\frac{1-\Phi(u_{n})}{1-\Phi(v_{n})}=\exp\!\left(-v_{n}\Delta_{n}-\frac{\Delta_{n}^{2}}{2}\right)\left(1+O\!\left(\frac{|\Delta_{n}|}{v_{n}}+\frac{1}{v_{n}^{2}}\right)\right).$ (19)
In particular, if $|\Delta_{n}|=O(v_{n}/n)$ and $v_{n}^{2}=O(\log n)$, then
$\frac{1-\Phi(u_{n})}{1-\Phi(v_{n})}=1+O\!\left(\frac{v_{n}^{2}}{n}+\frac{1}{v_{n}^{2}}\right),$ (20)
and hence
$1-\Phi(u_{n})=(1-\Phi(v_{n}))\left(1+O\!\left(\frac{1}{\log n}\right)\right).$ (21)
Proof of Lemma 1.
We use Result 2. For $x\to\infty$ write
$1-\Phi(x)=\frac{\varphi(x)}{x}\,m(x),\qquad m(x):=1-\frac{1}{x^{2}}+O\!\left(\frac{1}{x^{4}}\right).$
Since $m(x)=1+O(x^{-2})$ and $u_{n}\sim v_{n}\to\infty$, we have
$\frac{m(u_{n})}{m(v_{n})}=1+O\!\left(\frac{1}{v_{n}^{2}}\right),\qquad n\to\infty.$ (22)
Now take the ratio:
$\frac{1-\Phi(u_{n})}{1-\Phi(v_{n})}=\frac{\varphi(u_{n})}{\varphi(v_{n})}\cdot\frac{v_{n}}{u_{n}}\cdot\frac{m(u_{n})}{m(v_{n})}.$ (23)
First,
$\frac{\varphi(u_{n})}{\varphi(v_{n})}=\exp\!\left(-\frac{u_{n}^{2}-v_{n}^{2}}{2}\right)=\exp\!\left(-\frac{(v_{n}+\Delta_{n})^{2}-v_{n}^{2}}{2}\right)=\exp\!\left(-v_{n}\Delta_{n}-\frac{\Delta_{n}^{2}}{2}\right).$ (24)
Second,
$\frac{v_{n}}{u_{n}}=\frac{v_{n}}{v_{n}+\Delta_{n}}=\left(1+\frac{\Delta_{n}}{v_{n}}\right)^{-1}=1+O\!\left(\frac{|\Delta_{n}|}{v_{n}}\right),\qquad n\to\infty.$ (25)
Indeed, $u_{n}/v_{n}=1+\Delta_{n}/v_{n}\to 1$ since $\Delta_{n}=o(1)$ and $v_{n}\to\infty$, so $u_{n}\sim v_{n}$.
Substituting these three evaluations (22), (24), (25) into (23) yields (19).
Now, assume $|\Delta_{n}|=O(v_{n}/n)$ and $v_{n}^{2}=O(\log n)$. Then $|\Delta_{n}|/v_{n}=O(1/n)$ and $1/v_{n}^{2}=O(1/\log n)$, while
$v_{n}|\Delta_{n}|=O\!\left(\frac{v_{n}^{2}}{n}\right)=O\!\left(\frac{\log n}{n}\right),\,\,\,\Delta_{n}^{2}=O\!\left(\frac{v_{n}^{2}}{n^{2}}\right)=o\!\left(\frac{v_{n}^{2}}{n}\right).$
Therefore the exponential term satisfies
$\exp\!\left(-v_{n}\Delta_{n}-\frac{\Delta_{n}^{2}}{2}\right)=1+O(v_{n}|\Delta_{n}|)+O(\Delta_{n}^{2})=1+O\!\left(\frac{v_{n}^{2}}{n}\right),$
and multiplying by $1+O(|\Delta_{n}|/v_{n}+1/v_{n}^{2})$ we obtain
$\frac{1-\Phi(u_{n})}{1-\Phi(v_{n})}=1+O\!\left(\frac{v_{n}^{2}}{n}+\frac{1}{v_{n}^{2}}\right),$
which proves (20). Finally, under $v_{n}^{2}=O(\log n)$ we have $1/v_{n}^{2}=O(1/\log n)$ and $v_{n}^{2}/n=O(\log n/n)=o(1/\log n)$, hence
$\frac{1-\Phi(u_{n})}{1-\Phi(v_{n})}=1+O\!\left(\frac{1}{\log n}\right),$
which yields (21).
We will also need the following Cramér-type moderate deviation theorem.
Theorem 2 ((Petrov, 1975, Theorem 1, §VIII.2, p. 269; see also p. 323)).
Let $Y_{1},Y_{2},\ldots$ be i.i.d. random variables such that $\mathbb{E}Y_{1}=0$ and $\mathbb{E}Y_{1}^{2}=1$. Assume Cramér’s condition: there exists $H>0$ such that $\mathbb{E}e^{tY_{1}}<\infty$ for all $|t|<H$. Let $S_{n}:=\sum_{k=1}^{n}Y_{k}$ and let $x=x_{n}\geq 0$ satisfy $x=o(\sqrt{n})$. Then
$\frac{\mathbb{P}\!\left(S_{n}/\sqrt{n}>x\right)}{1-\Phi(x)}=\exp\!\left(\frac{x^{3}}{\sqrt{n}}\,\eta\!\left(\frac{x}{\sqrt{n}}\right)\right)\left(1+O\!\left(\frac{x+1}{\sqrt{n}}\right)\right),$
where $\eta(u)=\sum_{m\geq 0}a_{m}u^{m}$ is a power series whose coefficients depend on the cumulants of $Y_{1}$ and which converges for all sufficiently small $|u|$.
Lemma 2.
Fix $t\in\mathbb{R}$ and write $x_{n}:=x_{n}(t)$. Let $q_{n}(u)=\mathbb{P}(R_{1}>u)$, where
$R_{1}=\frac{1}{\sqrt{n-2}}\sum_{j=3}^{n}Y_{1j},\qquad Y_{ij}=\frac{X_{ij}-\mu}{\sigma}.$
Then the following hold.
(a) There exists a constant $C(t)<\infty$ such that, uniformly over all real sequences $\{u_{n}\}$ satisfying
$u_{n}=x_{n}+\Delta_{n},\qquad|\Delta_{n}|\leq\frac{C(t)}{\sqrt{n}},$ (26)
we have
$q_{n}(\alpha_{n}u_{n})=(1-\Phi(u_{n}))\left(1+O\!\left(\frac{(\log\log n)^{2}}{\log n}\right)\right),\qquad n\to\infty.$ (27)
(b) Under (26),
$q_{n}(\alpha_{n}(x_{n}-\Delta_{n}))-q_{n}(\alpha_{n}(x_{n}+\Delta_{n}))=2\Delta_{n}\,\varphi(x_{n})\left(1+O\!\left(\frac{(\log\log n)^{2}}{\log n}\right)\right),\qquad n\to\infty,$ (28)
uniformly in $|\Delta_{n}|\leq C(t)/\sqrt{n}$.
Proof of Lemma 2.
By definition, $R_{1}=\frac{1}{\sqrt{n-2}}\sum_{j=3}^{n}Y_{1j},$ where the summands are i.i.d., mean 0, variance $1$, and bounded. Hence Cramér’s condition holds, and Petrov’s theorem applies. Let m=n2m=n-2. Petrov’s Theorem 2 yields, uniformly for $u=o(\sqrt{m})$,
$\frac{\mathbb{P}(R_{1}>u)}{1-\Phi(u)}=\exp\!\left(\frac{u^{3}}{\sqrt{m}}\eta\!\left(\frac{u}{\sqrt{m}}\right)\right)\left(1+O\!\left(\frac{u+1}{\sqrt{m}}\right)\right),\qquad m\to\infty,$
where $\eta(\cdot)$ is analytic near 0.
Now in our regime $u=\alpha_{n}u_{n}$ with $u_{n}=x_{n}+\Delta_{n}$ and $|\Delta_{n}|\leq C/\sqrt{n}$, we have $u_{n}\sim x_{n}\sim\sqrt{2\log n}$ and $\alpha_{n}=1+O(1/n)$, hence
$u=\alpha_{n}u_{n}=O(\sqrt{\log n}),\qquad\frac{u}{\sqrt{m}}=O\!\left(\frac{\sqrt{\log n}}{\sqrt{n}}\right)\to 0.$ (29)
Thus $u=o(\sqrt{m})$ and Petrov Theorem 2 applies.
Moreover, from (29),
$\frac{u^{3}}{\sqrt{m}}=O\!\left(\frac{(\log n)^{3/2}}{\sqrt{n}}\right),\qquad\frac{u+1}{\sqrt{m}}=O\!\left(\frac{\sqrt{\log n}}{\sqrt{n}}\right).$
Both are $o((\log\log n)^{2}/\log n)$, because
$\frac{(\log n)^{3/2}}{\sqrt{n}}\Big/\frac{(\log\log n)^{2}}{\log n}=\frac{(\log n)^{5/2}}{\sqrt{n}(\log\log n)^{2}}\to 0.$
Hence,
$\mathbb{P}(R_{1}>u)=(1-\Phi(u))\left(1+O\!\left(\frac{(\log\log n)^{2}}{\log n}\right)\right).$ (30)
We now apply Lemma 1 with
$v_{n}=u_{n},\qquad u_{n}^{\star}:=\alpha_{n}u_{n},\qquad\Delta_{n}^{\star}=u_{n}^{\star}-v_{n}=(\alpha_{n}-1)u_{n}.$
We have $\alpha_{n}-1=O(1/n)$ and $u_{n}=O(\sqrt{\log n})$, so
$|\Delta_{n}^{\star}|=O\!\left(\frac{\sqrt{\log n}}{n}\right)=o(1).$
Also $v_{n}^{2}=u_{n}^{2}=O(\log n)$ and
$\frac{|\Delta_{n}^{\star}|}{v_{n}}=O\!\left(\frac{1}{n}\right),\qquad\frac{1}{v_{n}^{2}}=O\!\left(\frac{1}{\log n}\right).$
Therefore Lemma 1 gives
$\frac{1-\Phi(\alpha_{n}u_{n})}{1-\Phi(u_{n})}=1+O\!\left(\frac{1}{\log n}\right),\qquad n\to\infty,$ (31)
uniformly in $|\Delta_{n}|\leq C/\sqrt{n}$.
Combining (30) at $u=\alpha_{n}u_{n}$ with (31) yields
$q_{n}(\alpha_{n}u_{n})=(1-\Phi(\alpha_{n}u_{n}))\left(1+O\!\left(\frac{(\log\log n)^{2}}{\log n}\right)\right)=(1-\Phi(u_{n}))\left(1+O\!\left(\frac{(\log\log n)^{2}}{\log n}\right)\right),$
which is (27). This proves part (a).
Apply part (a) with $u_{n}=x_{n}\pm\Delta_{n}$ to get
$q_{n}(\alpha_{n}(x_{n}-\Delta_{n}))$ $=(1-\Phi(x_{n}-\Delta_{n}))\left(1+O\!\left(\frac{(\log\log n)^{2}}{\log n}\right)\right),$ (32)
$q_{n}(\alpha_{n}(x_{n}+\Delta_{n}))$ $=(1-\Phi(x_{n}+\Delta_{n}))\left(1+O\!\left(\frac{(\log\log n)^{2}}{\log n}\right)\right),$ (33)
uniformly for $|\Delta_{n}|\leq C/\sqrt{n}$.
Subtract (33) from (32):
$q_{n}(\alpha_{n}(x_{n}-\Delta_{n}))-q_{n}(\alpha_{n}(x_{n}+\Delta_{n}))$ $=\Bigl[(1-\Phi(x_{n}-\Delta_{n}))-(1-\Phi(x_{n}+\Delta_{n}))\Bigr]\left(1+O\!\left(\frac{(\log\log n)^{2}}{\log n}\right)\right).$ (34)
Now,
$(1-\Phi(x_{n}-\Delta_{n}))-(1-\Phi(x_{n}+\Delta_{n}))=\Phi(x_{n}+\Delta_{n})-\Phi(x_{n}-\Delta_{n})=\int_{x_{n}-\Delta_{n}}^{x_{n}+\Delta_{n}}\varphi(u)\,du.$ (35)
We estimate the integral by comparing $\varphi(u)$ to $\varphi(x_{n})$.
$\frac{\varphi(u)}{\varphi(x_{n})}=\exp\!\left(-\frac{u^{2}-x_{n}^{2}}{2}\right)=\exp\!\left(-(u-x_{n})x_{n}-\frac{(u-x_{n})^{2}}{2}\right).$
Since $|u-x_{n}|\leq|\Delta_{n}|\leq C/\sqrt{n}$ and $x_{n}\sim\sqrt{2\log n}$,
$|(u-x_{n})x_{n}|\leq\frac{C\sqrt{\log n}}{\sqrt{n}}\to 0,\qquad(u-x_{n})^{2}\leq\frac{C^{2}}{n}\to 0.$
Hence uniformly in the integration range,
$\varphi(u)=\varphi(x_{n})\left(1+O\!\left(\frac{\sqrt{\log n}}{\sqrt{n}}\right)\right).$ (36)
Substitute (36) into (35):
$\int_{x_{n}-\Delta_{n}}^{x_{n}+\Delta_{n}}\varphi(u)\,du=2\Delta_{n}\,\varphi(x_{n})\left(1+O\!\left(\frac{\sqrt{\log n}}{\sqrt{n}}\right)\right).$
Since $\sqrt{\log n}/\sqrt{n}=o((\log\log n)^{2}/\log n)$, and with (34) this yields (28). ∎
Lemma 3.
Fix $t\in\mathbb{R}$ and let $x_{n}:=x_{n}(t)$. Let $q_{n}(u)=\mathbb{P}(R_{1}>u)$, where
$R_{1}=\frac{1}{\sqrt{n-2}}\sum_{j=3}^{n}Y_{1j},\qquad Y_{ij}=\frac{X_{ij}-\mu}{\sigma}.$
Let
$\alpha_{n}:=\sqrt{\frac{n-1}{n-2}},\,\,\,\,A_{n}(y):=q_{n}\!\left(\alpha_{n}\left(x_{n}-\frac{y}{\sqrt{n-1}}\right)\right),\,\,\,\,B_{n}(y):=q_{n}\!\left(\alpha_{n}\left(x_{n}+\frac{y}{\sqrt{n-1}}\right)\right),$
and
$M_{n}(y):=\frac{A_{n}(y)+B_{n}(y)}{2}.$
Let $Y:=Y_{12}=(X_{12}-\mu)/\sigma$. Then there exists $C(t)<\infty$ such that
$\mathrm{Var}(M_{n}(Y))\leq C(t)\left(\frac{x_{n}\varphi(x_{n})}{n}\right)^{2},\qquad n\geq 3.$ (37)
In particular,
$\mathrm{Var}(M_{n}(Y))=O\!\left(\frac{(\log n)^{2}}{n^{4}}\right),\qquad n\to\infty.$ (38)
Proof of Lemma 3.
Fix $t\in\mathbb{R}$ and let $x_{n}=x_{n}(t)$. All implicit constants below may depend on $t$ but not on $n$.
Since $X_{12}\in[0,1]$ and $\sigma>0$ is fixed, $Y=(X_{12}-\mu)/\sigma$ is bounded: there exists $K<\infty$ such that $|Y|\leq K\quad\text{a.s.}$ Define
$\Delta:=\frac{Y}{\sqrt{n-1}},\qquad\text{so that}\qquad|\Delta|\leq\frac{K}{\sqrt{n-1}}\leq\frac{K}{\sqrt{n}}.$
Thus $|\Delta|\leq K/\sqrt{n}$, so Lemma 2(a) applies provided its window constant $C(t)$ is chosen $\geq K$ (which we may assume by enlarging $C(t)$ if needed).
Apply Lemma 2(a) with $u_{n}=x_{n}\pm\Delta$. There exists a deterministic sequence $\varepsilon_{n}=\varepsilon_{n}(t)$ such that
$\varepsilon_{n}=O\!\left(\frac{(\log\log n)^{2}}{\log n}\right),\qquad n\to\infty,$ (39)
and, uniformly for all realizations of $Y$ (equivalently, uniformly in $|\Delta|\leq K/\sqrt{n}$),
$A_{n}(Y)$ $=\bigl(1-\Phi(x_{n}-\Delta)\bigr)\,(1+\varepsilon_{n}),$
$B_{n}(Y)$ $=\bigl(1-\Phi(x_{n}+\Delta)\bigr)\,(1+\varepsilon_{n}).$
Consequently,
$M_{n}(Y)=(1+\varepsilon_{n})\,\widetilde{M}_{n}(\Delta),\qquad\widetilde{M}_{n}(\Delta):=\frac{(1-\Phi(x_{n}-\Delta))+(1-\Phi(x_{n}+\Delta))}{2}.$
Therefore,
$\mathrm{Var}(M_{n}(Y))=(1+\varepsilon_{n})^{2}\,\mathrm{Var}\bigl(\widetilde{M}_{n}(\Delta)\bigr).$ (40)
Therefore it suffices to bound $\mathrm{Var}(\widetilde{M}_{n}(\Delta))$.
Define $g(u):=1-\Phi(u)$. Then $g^{\prime}(u)=-\varphi(u)$ and $g^{\prime\prime}(u)=-\varphi^{\prime}(u)$. Taylor’s theorem gives: for each $\Delta$, there exist $\theta_{1},\theta_{2}\in(-1,1)$ such that
$g(x_{n}-\Delta)$ $=g(x_{n})-\Delta g^{\prime}(x_{n})+\frac{\Delta^{2}}{2}g^{\prime\prime}(x_{n}-\theta_{1}\Delta)=g(x_{n})+\Delta\varphi(x_{n})-\frac{\Delta^{2}}{2}\varphi^{\prime}(x_{n}-\theta_{1}\Delta),$ (41)
$g(x_{n}+\Delta)$ $=g(x_{n})+\Delta g^{\prime}(x_{n})+\frac{\Delta^{2}}{2}g^{\prime\prime}(x_{n}+\theta_{2}\Delta)=g(x_{n})-\Delta\varphi(x_{n})-\frac{\Delta^{2}}{2}\varphi^{\prime}(x_{n}+\theta_{2}\Delta).$ (42)
Averaging (41) and (42) yields:
$\widetilde{M}_{n}(\Delta)=g(x_{n})-\frac{\Delta^{2}}{4}\Bigl(\varphi^{\prime}(x_{n}-\theta_{1}\Delta)+\varphi^{\prime}(x_{n}+\theta_{2}\Delta)\Bigr).$ (43)
Therefore,
$\mathrm{Var}\bigl(\widetilde{M}_{n}(\Delta)\bigr)=\mathrm{Var}\!\left(\frac{\Delta^{2}}{4}\Bigl(\varphi^{\prime}(x_{n}-\theta_{1}\Delta)+\varphi^{\prime}(x_{n}+\theta_{2}\Delta)\Bigr)\right).$
Recall $\varphi^{\prime}(u)=-u\varphi(u)$. We claim that for all sufficiently large $n$ (depending on $t$), there exists $C<\infty$ such that
$\sup_{|\theta|\leq 1}\ \sup_{|\Delta|\leq K/\sqrt{n}}\ \bigl|\varphi^{\prime}(x_{n}+\theta\Delta)\bigr|\leq C\,x_{n}\varphi(x_{n}).$ (44)
Indeed, for such $\theta,\Delta$,
$\bigl|\varphi^{\prime}(x_{n}+\theta\Delta)\bigr|=|x_{n}+\theta\Delta|\,\varphi(x_{n}+\theta\Delta)\leq(|x_{n}|+|\Delta|)\,\varphi(x_{n}+\theta\Delta).$
Moreover,
$\frac{\varphi(x_{n}+\theta\Delta)}{\varphi(x_{n})}=\exp\!\left(-\theta x_{n}\Delta-\frac{\theta^{2}\Delta^{2}}{2}\right).$
Since $|x_{n}|=O(\sqrt{\log n})$ and $|\Delta|\leq K/\sqrt{n}$, we have $|x_{n}\Delta|\leq K\sqrt{\log n}/\sqrt{n}\to 0$ and $\Delta^{2}\leq K^{2}/n\to 0$, so this ratio is bounded by positive constant for all large $n$. Also $|\Delta|=o(x_{n})$, hence $|x_{n}|+|\Delta|\leq 2x_{n}$ for all large $n$. Combining these bounds yields (44).
Let
$H_{n}(\Delta):=\frac{1}{4}\Bigl(\varphi^{\prime}(x_{n}-\theta_{1}\Delta)+\varphi^{\prime}(x_{n}+\theta_{2}\Delta)\Bigr).$
By (44), for all large $n$ and all admissible $\Delta$,
$|H_{n}(\Delta)|\leq C\,x_{n}\varphi(x_{n}).$ (45)
Then (43) rewrites as
$\widetilde{M}_{n}(\Delta)=g(x_{n})-\Delta^{2}\,H_{n}(\Delta),$
hence,
$\mathrm{Var}\bigl(\widetilde{M}_{n}(\Delta)\bigr)=\mathrm{Var}\bigl(\Delta^{2}H_{n}(\Delta)\bigr)\leq\mathbb{E}\bigl[(\Delta^{2}H_{n}(\Delta))^{2}\bigr].$
Using (45),
$\mathrm{Var}\bigl(\widetilde{M}_{n}(\Delta)\bigr)\leq C\,x_{n}^{2}\varphi(x_{n})^{2}\,\mathbb{E}[\Delta^{4}].$ (46)
Now $\Delta=Y/\sqrt{n-1}$, so
$\mathbb{E}[\Delta^{4}]=\frac{\mathbb{E}[Y^{4}]}{(n-1)^{2}}\leq\frac{C}{n^{2}},$ (47)
since $Y$ is bounded. Substituting (47) into (46) yields
$\mathrm{Var}\bigl(\widetilde{M}_{n}(\Delta)\bigr)\leq C\left(\frac{x_{n}\varphi(x_{n})}{n}\right)^{2}.$ (48)
Insert (48) into (40):
$\mathrm{Var}(M_{n}(Y))\leq(1+\varepsilon_{n})^{2}\,C\left(\frac{x_{n}\varphi(x_{n})}{n}\right)^{2}.$
Since $\varepsilon_{n}\to 0$ by (39), $(1+\varepsilon_{n})^{2}$ is bounded, and we obtain (37).
From (53), we have
$\varphi(x_{n})=\frac{e^{-t}\sqrt{2\log n}}{n}\left(1+O\!\left(\frac{(\log\log n)^{2}}{\log n}\right)\right),$
and since $x_{n}=O(\sqrt{\log n})$ we have $x_{n}\varphi(x_{n})=O(\log n/n)$. Substituting into (37) gives (38). ∎
Appendix B Proof of Proposition 1
Proof of Proposition 1.
Fix $t\in\mathbb{R}$.
Recall from (10) that $p_{n}(t)=\mathbb{P}\!\left(s_{1}^{\star}(n)>x_{n}(t)\right),\,\,\,\lambda_{n}(t)=n\,p_{n}(t)$ and $s_{1}(n)=\sum_{j=2}^{n}X_{1j}.$ Define
$Y_{j}:=\frac{X_{1,j+1}-\mu}{\sigma},\qquad j=1,2,\ldots.$
Then $\{Y_{j}\}$ are i.i.d., $\mathbb{E}Y_{1}=0$, $Var(Y_{1})=1$, and
$s_{1}^{\star}(n)=\frac{s_{1}(n)-(n-1)\mu}{\sqrt{n-1}\,\sigma}=\frac{1}{\sqrt{n-1}}\sum_{j=1}^{n-1}Y_{j}.$
Therefore,
$p_{n}(t)=\mathbb{P}\!\left(\frac{1}{\sqrt{n-1}}\sum_{j=1}^{n-1}Y_{j}>x_{n}(t)\right).$ (49)
Moreover, since $X_{ij}\in D\subset[0,1]$ is bounded, the Cramér’s condition in Theorem 2 holds.
From (8) we have $x_{n}(t)\sim\sqrt{2\log n}$ as $n\to\infty$, hence $x_{n}(t)=o(\sqrt{n-1})\qquad(n\to\infty).$ Thus Theorem 2 is applicable. Applying Theorem 2 to (49) yields
$\frac{p_{n}(t)}{1-\Phi(x_{n}(t))}=\exp\!\left(\frac{x_{n}(t)^{3}}{\sqrt{n-1}}\,\eta\!\left(\frac{x_{n}(t)}{\sqrt{n-1}}\right)\right)\left(1+O\!\left(\frac{x_{n}(t)+1}{\sqrt{n-1}}\right)\right),\qquad n\to\infty,$
where $\eta(\cdot)$ is analytic near 0 (hence bounded in a neighborhood of 0). Since $x_{n}(t)=O(\sqrt{\log n})$, we have
$\frac{x_{n}(t)^{3}}{\sqrt{n-1}}=O\!\left(\frac{(\log n)^{3/2}}{\sqrt{n}}\right)\to 0,\qquad\frac{x_{n}(t)+1}{\sqrt{n-1}}=O\!\left(\frac{\sqrt{\log n}}{\sqrt{n}}\right)\to 0,$
and hence
$\frac{p_{n}(t)}{1-\Phi(x_{n}(t))}=1+O\!\left(\frac{(\log n)^{3/2}}{\sqrt{n}}\right),\qquad n\to\infty.$ (50)
Note that $(\log n)^{3/2}/\sqrt{n}=o((\log\log n)^{2}/\log n)$.
By Result 2,
$1-\Phi(x)=\frac{\varphi(x)}{x}\left(1+O\!\left(\frac{1}{x^{2}}\right)\right),\qquad x\to\infty.$ (51)
Since $x_{n}(t)\sim\sqrt{2\log n}\to\infty$, (51) applies with $x=x_{n}(t)$:
$1-\Phi(x_{n}(t))=\frac{\varphi(x_{n}(t))}{x_{n}(t)}\left(1+O\!\left(\frac{1}{x_{n}(t)^{2}}\right)\right)=\frac{\varphi(x_{n}(t))}{x_{n}(t)}\left(1+O\!\left(\frac{1}{\log n}\right)\right).$ (52)
Recall (7) and define $u_{n}:=\sqrt{2\log n}$ and $c_{n}:=\log\log n+\log(4\pi)$, so that $b_{n}=u_{n}-\frac{c_{n}}{2u_{n}},\,\,\,a_{n}=\frac{1}{u_{n}}.$ A direct expansion gives
$x_{n}(t)^{2}$ $=(b_{n}+a_{n}t)^{2}=b_{n}^{2}+2a_{n}b_{n}\,t+a_{n}^{2}t^{2}$
$=\left(u_{n}-\frac{c_{n}}{2u_{n}}\right)^{2}+2\left(\frac{1}{u_{n}}\right)\left(u_{n}-\frac{c_{n}}{2u_{n}}\right)t+O\!\left(\frac{1}{\log n}\right)$
$=u_{n}^{2}-c_{n}+2t+O\!\left(\frac{(\log\log n)^{2}}{\log n}\right)\qquad(n\to\infty),$
since $u_{n}^{2}=2\log n$ and $c_{n}^{2}/u_{n}^{2}=O((\log\log n)^{2}/\log n)$. Therefore
$\varphi(x_{n}(t))=(2\pi)^{-1/2}\exp\!\left(-\frac{x_{n}(t)^{2}}{2}\right)=(2\pi)^{-1/2}\exp\!\left(-\log n+\frac{c_{n}}{2}-t+O\!\left(\frac{(\log\log n)^{2}}{\log n}\right)\right).$
Using $\exp(c_{n}/2)=\exp((\log\log n)/2)\exp((\log(4\pi))/2)=\sqrt{\log n}\,\sqrt{4\pi}$, we obtain
$\varphi(x_{n}(t))=\frac{e^{-t}\sqrt{2\log n}}{n}\left(1+O\!\left(\frac{(\log\log n)^{2}}{\log n}\right)\right).$ (53)
Substitute (53) into (52) and use $x_{n}(t)\sim\sqrt{2\log n}$, we obtain
$1-\Phi(x_{n}(t))=\frac{e^{-t}}{n}\left(1+O\!\left(\frac{(\log\log n)^{2}}{\log n}\right)\right).$ (54)
From (50) and (54),
$p_{n}(t)=\left(1-\Phi(x_{n}(t))\right)\left(1+O\!\left(\frac{(\log n)^{3/2}}{\sqrt{n}}\right)\right)=\frac{e^{-t}}{n}\left(1+O\!\left(\frac{(\log\log n)^{2}}{\log n}\right)\right),$
and therefore
$\lambda_{n}(t)=n\,p_{n}(t)=e^{-t}\left(1+O\!\left(\frac{(\log\log n)^{2}}{\log n}\right)\right),\qquad n\to\infty.$
Appendix C Proof of Proposition 2
Proof of Proposition 2.
Fix $t\in\mathbb{R}$ and let $x_{n}:=x_{n}(t).$ Recall that
$I_{i}^{(n)}(t)=\mathbf{1}\{s_{i}^{\star}(n)>x_{n}\},\qquad p_{n}(t)=\mathbb{P}(I_{1}^{(n)}(t)=1),\qquad\lambda_{n}(t)=np_{n}(t).$
We shall compute
$\mathrm{Cov}\!\left(I_{1}^{(n)}(t),I_{2}^{(n)}(t)\right)=\mathbb{P}(I_{1}^{(n)}(t)=1,I_{2}^{(n)}(t)=1)-p_{n}(t)^{2}.$
Define
$Y_{ij}:=\frac{X_{ij}-\mu}{\sigma},\qquad i\neq j.$
Under Model $M_{[0,1]}$, we have $\mu=\mathbb{E}(X_{ij})=1/2$ by (6). Also, $X_{ij}\in[0,1]$ implies $|Y_{ij}|\leq 1/\sigma$ is bounded.
For player $1$,
$s_{1}^{\star}(n)=\frac{1}{\sqrt{n-1}}\sum_{j=2}^{n}Y_{1j}=\frac{1}{\sqrt{n-1}}\Bigl(Y_{12}+\sum_{j=3}^{n}Y_{1j}\Bigr).$
For player $2$,
$s_{2}^{\star}(n)=\frac{1}{\sqrt{n-1}}\sum_{j\neq 2}Y_{2j}=\frac{1}{\sqrt{n-1}}\Bigl(Y_{21}+\sum_{j=3}^{n}Y_{2j}\Bigr).$
Using $X_{21}=1-X_{12}$ and $\mu=1/2$, we get
$Y_{21}=\frac{X_{21}-1/2}{\sigma}=\frac{(1-X_{12})-1/2}{\sigma}=-\frac{X_{12}-1/2}{\sigma}=-Y_{12}.$
Setting, $Y:=Y_{12},$ we obtain
$s_{1}^{\star}(n)=\frac{1}{\sqrt{n-1}}\Bigl(Y+\sum_{j=3}^{n}Y_{1j}\Bigr),\qquad s_{2}^{\star}(n)=\frac{1}{\sqrt{n-1}}\Bigl(-Y+\sum_{j=3}^{n}Y_{2j}\Bigr).$ (55)
Now define:
$R_{1}:=\frac{1}{\sqrt{n-2}}\sum_{j=3}^{n}Y_{1j},\qquad R_{2}:=\frac{1}{\sqrt{n-2}}\sum_{j=3}^{n}Y_{2j}.$
Then (55) becomes
$s_{1}^{\star}(n)=\sqrt{\frac{n-2}{n-1}}\,R_{1}+\frac{Y}{\sqrt{n-1}},\qquad s_{2}^{\star}(n)=\sqrt{\frac{n-2}{n-1}}\,R_{2}-\frac{Y}{\sqrt{n-1}}.$ (56)
The collection of match vectors $\{(X_{ij},X_{ji}):1\leq i<j\leq n\}$ is independent by assumption. Therefore:
  • $Y=Y_{12}$ depends only on the match $(1,2)$;
  • $\{Y_{1j}:j=3,\dots,n\}$ depends only on matches $(1,j)$, $j\geq 3$;
  • $\{Y_{2j}:j=3,\dots,n\}$ depends only on matches $(2,j)$, $j\geq 3$;
and these three families correspond to disjoint sets of matches, hence are independent. Consequently,
$Y,\ R_{1},\ R_{2}\text{ are independent, and }R_{1}\stackrel{{\scriptstyle d}}{{=}}R_{2}.$ (57)
Introduce the notations
$\alpha_{n}:=\sqrt{\frac{n-1}{n-2}},\qquad h_{n}(y):=\frac{y}{\sqrt{n-1}}.$
Then from (56),
$\{s_{1}^{\star}(n)>x_{n}\}$ $=\left\{R_{1}>\alpha_{n}\bigl(x_{n}-h_{n}(Y)\bigr)\right\},$ (58)
$\{s_{2}^{\star}(n)>x_{n}\}$ $=\left\{R_{2}>\alpha_{n}\bigl(x_{n}+h_{n}(Y)\bigr)\right\}.$ (59)
Define
$q_{n}(u):=\mathbb{P}(R_{1}>u),\qquad u\in\mathbb{R}.$
Conditioning on $Y$ and using (57)–(59),
$\mathbb{P}(I_{1}^{(n)}(t)=1,I_{2}^{(n)}(t)=1\mid Y)$ $=\mathbb{P}\!\left(R_{1}>\alpha_{n}(x_{n}-h_{n}(Y))\mid Y\right)\mathbb{P}\!\left(R_{2}>\alpha_{n}(x_{n}+h_{n}(Y))\mid Y\right)$
$=q_{n}\!\left(\alpha_{n}(x_{n}-h_{n}(Y))\right)\,q_{n}\!\left(\alpha_{n}(x_{n}+h_{n}(Y))\right).$
Taking expectation,
$\mathbb{P}(I_{1}^{(n)}(t)=1,I_{2}^{(n)}(t)=1)=\mathbb{E}\!\left[q_{n}\!\left(\alpha_{n}(x_{n}-h_{n}(Y))\right)\,q_{n}\!\left(\alpha_{n}(x_{n}+h_{n}(Y))\right)\right].$ (60)
Similarly,
$p_{n}(t)=\mathbb{P}(I_{1}^{(n)}(t)=1)=\mathbb{E}\!\left[q_{n}\!\left(\alpha_{n}(x_{n}-h_{n}(Y))\right)\right].$ (61)
Therefore, everything reduces to obtaining sufficiently accurate uniform asymptotics for $q_{n}(\alpha_{n}(x_{n}\pm h_{n}(Y)))$ when $h_{n}(Y)=O(n^{-1/2})$.
Let
$A_{n}(y):=q_{n}\!\left(\alpha_{n}\left(x_{n}-\frac{y}{\sqrt{n-1}}\right)\right),\,\,\,\,\,B_{n}(y):=q_{n}\!\left(\alpha_{n}\left(x_{n}+\frac{y}{\sqrt{n-1}}\right)\right).$ (62)
From (60), (61) and (62), we obtain
$\mathbb{P}(I_{1}^{(n)}(t)=1,I_{2}^{(n)}(t)=1)$ $=\mathbb{E}\!\left[A_{n}(Y)\,B_{n}(Y)\right],\,\,\,p_{n}(t)=\mathbb{E}[A_{n}(Y)].$ (63)
Define
$M_{n}(Y):=\frac{A_{n}(Y)+B_{n}(Y)}{2},\qquad\Delta_{n}(Y):=\frac{A_{n}(Y)-B_{n}(Y)}{2}.$ (64)
Then
$A_{n}(Y)=M_{n}(Y)+\Delta_{n}(Y),\qquad B_{n}(Y)=M_{n}(Y)-\Delta_{n}(Y).$
Recall that $Y=\frac{X_{12}-\mu}{\sigma},\,\,\,\,\mu=\tfrac{1}{2}.$ Since the distribution of $X_{12}$ is symmetric about $1/2$, we have $Y\stackrel{{\scriptstyle d}}{{=}}-Y.$
Moreover, by definition,
$\Delta_{n}(y)=\frac{1}{2}\left[q_{n}\!\left(\alpha_{n}\left(x_{n}-\frac{y}{\sqrt{n-1}}\right)\right)-q_{n}\!\left(\alpha_{n}\left(x_{n}+\frac{y}{\sqrt{n-1}}\right)\right)\right],$
which implies that $\Delta_{n}(\cdot)$ is an odd function, i.e.,
$\Delta_{n}(-y)=-\Delta_{n}(y).$
Therefore,
$\mathbb{E}[\Delta_{n}(Y)]=\mathbb{E}[\Delta_{n}(-Y)]=-\mathbb{E}[\Delta_{n}(Y)],$
and hence
$\mathbb{E}[\Delta_{n}(Y)]=0.$
Hence
$A_{n}(Y)B_{n}(Y)=M_{n}(Y)^{2}-\Delta_{n}(Y)^{2}.$ (65)
By (63) and (65),
$\mathrm{Cov}(I_{1}^{(n)}(t),I_{2}^{(n)}(t))$ $=\mathbb{E}[A_{n}(Y)B_{n}(Y)]-\bigl(\mathbb{E}[A_{n}(Y)]\bigr)^{2}$
$=\Bigl(\mathbb{E}[M_{n}(Y)^{2}]-\bigl(\mathbb{E}[M_{n}(Y)]\bigr)^{2}\Bigr)-\mathbb{E}[\Delta_{n}(Y)^{2}]$
$=\mathrm{Var}(M_{n}(Y))-\mathbb{E}[\Delta_{n}(Y)^{2}].$ (66)
By (64),
$2\Delta_{n}(Y)=A_{n}(Y)-B_{n}(Y)=q_{n}\!\left(\alpha_{n}\left(x_{n}-\frac{Y}{\sqrt{n-1}}\right)\right)-q_{n}\!\left(\alpha_{n}\left(x_{n}+\frac{Y}{\sqrt{n-1}}\right)\right).$
Since $Y$ is bounded, $|Y|\leq K$ a.s., hence $|\Delta_{n}|=|Y|/\sqrt{n-1}\leq K/\sqrt{n}$.
Thus Lemma 2(b) applies after choosing its window constant $C(t)\geq K$:
$2\Delta_{n}(Y)=2\cdot\frac{Y}{\sqrt{n-1}}\,\varphi(x_{n})\left(1+O\!\left(\frac{(\log\log n)^{2}}{\log n}\right)\right).$
Therefore
$\Delta_{n}(Y)=\frac{Y}{\sqrt{n-1}}\,\varphi(x_{n})\left(1+O\!\left(\frac{(\log\log n)^{2}}{\log n}\right)\right),$
and
$\Delta_{n}(Y)^{2}=\frac{Y^{2}}{n-1}\,\varphi(x_{n})^{2}\left(1+O\!\left(\frac{(\log\log n)^{2}}{\log n}\right)\right),$
where we used that $(1+O(\varepsilon_{n}))^{2}=1+O(\varepsilon_{n})$. Taking expectations and using $\mathbb{E}[Y^{2}]=1$,
$\mathbb{E}[\Delta_{n}(Y)^{2}]=\frac{\varphi(x_{n})^{2}}{n-1}\left(1+O\!\left(\frac{(\log\log n)^{2}}{\log n}\right)\right).$ (67)
From (53),
$\varphi(x_{n})=\frac{e^{-t}\sqrt{2\log n}}{n}\left(1+O\!\left(\frac{(\log\log n)^{2}}{\log n}\right)\right),$
so
$\varphi(x_{n})^{2}=\frac{2e^{-2t}\log n}{n^{2}}\left(1+O\!\left(\frac{(\log\log n)^{2}}{\log n}\right)\right).$ (68)
Substitute (68) into (67):
$\mathbb{E}[\Delta_{n}(Y)^{2}]=\frac{2e^{-2t}\log n}{n^{3}}\left(1+O\!\left(\frac{(\log\log n)^{2}}{\log n}\right)\right).$ (69)
By Lemma 3,
$\mathrm{Var}(M_{n}(Y))=O\!\left(\frac{(\log n)^{2}}{n^{4}}\right).$ (70)
Insert (69) and (70) into (66):
$\mathrm{Cov}(I_{1}^{(n)}(t),I_{2}^{(n)}(t))$ $=\mathrm{Var}(M_{n}(Y))-\mathbb{E}[\Delta_{n}(Y)^{2}]$
$=-\frac{2e^{-2t}\log n}{n^{3}}\left(1+O\!\left(\frac{(\log\log n)^{2}}{\log n}\right)\right)+O\!\left(\frac{(\log n)^{2}}{n^{4}}\right).$
Since $O((\log n)^{2}/n^{4})=\frac{2e^{-2t}\log n}{n^{3}}\cdot O(\log n/n)$, and $\log n/n=o((\log\log n)^{2}/\log n)$, we obtain,
$\mathrm{Cov}\!\left(I_{1}^{(n)}(t),I_{2}^{(n)}(t)\right)=-\frac{2\,e^{-2t}\log n}{n^{3}}\left(1+O\!\left(\frac{(\log\log n)^{2}}{\log n}\right)\right),\qquad n\to\infty.$
Acknowledgement
This research is supported in part by BSF grant 2020063.